OpenStudy (anonymous):

Multiply and simplify: (x²+2x+2)(x²-2x+2) Please explain how you did it! Thank you!

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

nobody helped me.

OpenStudy (anonymous):

answer is \[x^4+4\] and you have a choice. you can do 3 times 3 = 9 multiplications

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

you have to do 9 multiplications, but what is more interesting is to show \[x^4+4=(x^2+2x+2)(x^2-2x+2)\]

OpenStudy (anonymous):

wait why do you have to do nine? Not only three?

OpenStudy (anonymous):

ok suppose you wanted to multiply two 3-digit numbers together. how many multiplications would you have to do? say \[213\times 431\]

OpenStudy (anonymous):

WEll i understand now!

OpenStudy (anonymous):

really? great!

OpenStudy (anonymous):

Yeah except i still need help! Just not on that part!

OpenStudy (anonymous):

now in numbers we do it from the bottom up, starting at the ones place, then tens and so on. but with polynomials we usually to it from the top down because there is no borrowing or carrying

OpenStudy (anonymous):

what do we do?

OpenStudy (anonymous):

so if you just want \[(x^2-2x+2)(x^2+2x+2)\] you have to do the following nine multiplications \[x^2(x^2+2x+2)-2x(x^2+2x+2)+2(x^2+2x+2)\] \[x^4+3x^3+2x^2-2x^3-4x^2-4x+2x^2+4x+4\]

OpenStudy (anonymous):

see we have nine terms! nine, because of 3 terms times 3 terms means nine multiplications

OpenStudy (anonymous):

Yeah I get it

OpenStudy (anonymous):

now when you collect terms you get \[x^4+4\] because in this case a bunch of stuff adds up to zero

OpenStudy (anonymous):

Ok thanks so much!